OpenStudy (anonymous):

Let f(x) = sin(x) + icos(x). Show that |f(x)| = 1. I'm not even sure where to start here...

OpenStudy (anonymous):

hey its too much easy... |f(x)|=sqrt((1)^2+(1)^2)=sqrt(2)

OpenStudy (anonymous):

here we will take its real part 1 will cos(x) and also 1 from imagnary part i(sinx) hey u have written wrong equation.. its cos(x)-isin(x)

OpenStudy (anonymous):

The question is right. I know DeMoivres, but in this case it's sin(x) + icos(x).

OpenStudy (anonymous):

ah oookkk ya its absolute value

OpenStudy (anonymous):

So far I have LHS = \[\frac{ e ^{ix}-e^{-ix} }{ 2 } + \frac{ e ^{ix}+e^{-ix} }{ 2 }i\] Have I completely overthought the problem ? :)

OpenStudy (anonymous):

no i is denominator from demorvie`s, but in this case I did`nt get u how u got this result

OpenStudy (anonymous):

ok am gonna solve its prove and implement according to this condition the `ll msg u here ok

OpenStudy (anonymous):

Thanks :) I'm so confused...

OpenStudy (anonymous):

hahahhah i just wanna say write i in denominator ok i mean (eix−e−ix)2+(eix+e−ix)2i

OpenStudy (anonymous):

Hmm... I don't get it. How does that show that |sin(x) + icos(x)| = 1?

OpenStudy (anonymous):

oh i see wait i solve it and post the diagrame okk buddy

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

daer i can`t .. i have seen the same result for it sqareroot 2 iz answer for both condition.. cos when we see cosx+isinx,, or sinx+icosx... taking coifficients we get the same result..

OpenStudy (anonymous):

i thinks its wrong question or some thing missing ... can u give the link where did u got it from ???

OpenStudy (anonymous):

Yep, it's question 1 iv) from this paper: http://www.scribd.com/doc/112148611/MATH1131-Mathematics-1A-1 Thanks :)

OpenStudy (anonymous):

pdf attached here.

OpenStudy (anonymous):

u got ur answer or not /??

OpenStudy (anonymous):

Nope, heh. Not yet :(

OpenStudy (anonymous):

ok if its difficault then msg me ur question i will solve it my self and will post the pic here ok

OpenStudy (anonymous):

Yeah I don't know how to get the function to = 1...

OpenStudy (accessdenied):

I am thinking that you are using the fact that |f(x)| represents the exact distance between it and 0 on the complex plane... Notice that you can create a right triangle with the real part of f(x) and the imaginary part of x on a complex plane:  So, here, Re(f) = sin(x) and Im(f) = cos(x). We drop off the i since we're dealing with distances only for the imaginary part. Using Pythagorean theorem, |f(x)|^2 = (Re(f))^2 + (Im(f))^2 = sin^2 x + cos^2 x. Since we have a neat trigonometric identity here, we can use it to reduce sin^2 x + cos^2 x as 1. Take the square root of both sides and we're done. :)